In order to buy an automated packing machine for your factory, the cost would be $150,000. You have inquired at the bank about borrowing that amount and the bank would charge you a rate of 9.5% on the loan. The packing machine would be used for 10 years, at the end of which it would have no salvage value. It falls in an asset class with an allowable CCA rate of 10% per year. Use of the machine would reduce before-tax costs by $30,000 per year. Use of the machine would reduce packing machine from the manufacturer for lease payments of $25,000 per year (due at the beginning of each year). your firm's tax rate is 35%. Assuming that you will obtrain the new packing machine in some way, should you lease or buy it?
Answer #1

Buying option:

The loan amount is repayable together with interest of 9.5% on loan amount and is repayable over a period of 10 years assuming at the end of each year.

PVAF at the rate of 9.5% for 10 years is 6.2788

Annual payment= Cost if the machine ÷ PVAF = 150000÷6.2788= $23890 (rounded off)

Debt repayment schedule

End of

Year

Total payment

$

Interest

$

Principal

$

Principal amount

Outstanding $

1 23890

14250

(150000*9.5%)

9640

(23890-14250)

140360

(150000-9640)

2 23890

13334

(140360*9.5%)

10556

(23890-10556)

129804

(140360-10556)

3 23890 12331 11559 118245
4 23890 11233 12657 105588
5 23890 10031 13859 91729
6 23890 8714 15176 76553
7 23890 7273 16617 59936
8 23890 5694 18196 41740
9 23890 3965 19925 21815
10 23890

2075

(Rounded off)

21815 ------
Schedule of Cash Outflows ($)
1 2 3 4 3+4

5

6 7 6*7
Year Debt payment Interest Dep

Tax shield

(3+4)*0.35

Cash outflows

2-5

PV

factor@6.175%

PV
1 23890 14250 15000 29250 10238 13652 0.9418 12857
2 23890 13334 15000 28334 9917 13973 0.8871 12395
3 23890 12331 15000 27331 9566 14324 0.8355 11968
4 23890 11233 15000 26233 9182 14708 0.7869 11574
5 23890 10031 15000 25031 8761 15129 0.7411 11212
6 23890 8714 15000 23714 8300 15590 0.6980 10882
7 23890 7273 15000 22273 7796 16094 0.6574 10580
8 23890 5694 15000 20694 7243 16647 0.6192 10308
9 23890 3965 15000 18965 6638 17252 0.5832 10061
10 23890 2075 15000 17075 5976 17914 0.5493 9840
111677

There is also saving of $30000 per year before tax

Saving after tax=30000*(1-0.35)=$19500

PV of saving = 19500*PVAF@6.175%. for 10 years

PV of saving= 19500*7.2994= $142338

Net Saving for buying option = 142338-111677= $30661

Lease option:

Lease rent for years =$25000

Lease rental for Y1 beginning to Y10 beginning = 25000-tax saving

=25000*(1-0.35)= $16250

PV of outflows =25000+16250*PVAF@6.175% for 9years

=25000+16250*6.750= $134688

Since the PV of outflows is lower in the buying option it is preferable to avail of the loan and purchase the machine.
Answer #2

To determine whether you should lease or buy the packing machine, we need to compare the after-tax costs of each option over the 10-year period. Let's calculate the after-tax costs for both scenarios:

Leasing: Lease payments per year: $25,000 Tax rate: 35%

After-tax cost of leasing per year = Lease payments - Tax savings on lease payments = $25,000 - ($25,000 * 35%) = $25,000 - $8,750 = $16,250

Total after-tax cost of leasing over 10 years = After-tax cost of leasing per year * Number of years = $16,250 * 10 = $162,500

Buying: Cost of the packing machine: $150,000 CCA rate: 10% Tax rate: 35% Annual tax savings due to CCA = Cost of the packing machine * CCA rate * Tax rate = $150,000 * 10% * 35% = $5,250

Net after-tax cost of buying per year = Before-tax cost savings - Annual tax savings due to CCA = $30,000 - $5,250 = $24,750

Total after-tax cost of buying over 10 years = Net after-tax cost of buying per year * Number of years = $24,750 * 10 = $247,500

Comparing the total after-tax costs: Leasing: $162,500 Buying: $247,500

Based on the comparison, leasing the packing machine would result in a lower after-tax cost over the 10-year period. Therefore, it would be more cost-effective to lease the machine rather than buying it.

answered by: Hydra Master